OpenStudy (anonymous):

Simplify. -(5/8)(5x)(8/5)(12y)(0)

OpenStudy (anonymous):

these are fractions

OpenStudy (anonymous):

A. –60xy B. –60 C. 0 D. 60xy

Nnesha (nnesha):

you can do it in one sec if you know 0 times something = what

OpenStudy (anonymous):

oh right lol thx didn't see that

OpenStudy (anonymous):

the answer is C

Nnesha (nnesha):

\[-(\frac{ 5 }{ 8})(5x)(\frac{ 8 }{ 5 })(12y)\] let's say there is no 0 then multiply the coefficient \[-\frac{ 5 }{ 8 } \times \frac{ 8 }{ 5 }\] 5/8 and 8/5 are reciprocal

Nnesha (nnesha):

so you can cancel them \[-\frac{ \cancel{5} }{ 8 } \times \frac{ 8 }{ \cancel{5} }\] and 8 also cancels left with -1

OpenStudy (anonymous):

how do u do that?

Nnesha (nnesha):

-1 (5x)(12y)

Nnesha (nnesha):

that's arleady in the question `-(5/8)`(5x) `(8/5)`(12y)(0) when you multiply reciprocal you will get one

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

how do u write all those bold letter fractions and stuff

Nnesha (nnesha):

np this little steps help you to solve question quickly in math

Nnesha (nnesha):

i use equation tool :=)

OpenStudy (anonymous):

oh ok thx

Nnesha (nnesha):

http://prntscr.com/8opbtl